π‘₯2 + 𝑦2 βˆ’ 8π‘₯ + 2𝑦 βˆ’ 32 = 0?

1
Expert's answer
2021-11-17T17:16:54-0500

Since the question is incomplete, we find the type of conic section and its parameter.

Solution:

"𝑥^2 + 𝑦^2 − 8𝑥 +2𝑦 − 32 = 0
\\ \Rightarrow x^2-8x+y^2+2y=32
\\ \Rightarrow (x-4)^2-4^2+(y+1)^2-1^2=32
\\ \Rightarrow (x-4)^2+(y+1)^2=32+17
\\ \Rightarrow (x-4)^2+(y+1)^2=49
\\\Rightarrow (x-4)^2+(y+1)^2=7^2"

It is a circle. On comparing with "(x-h)^2+(y-k)^2=r^2", we get,

Centre"=(h,k)=(4,-1)" and radius"=r=7" units.